OpenStudy (anonymous):

help? D: What is the simplified form of the quantity 4 x squared minus 25 over the quantity 2 x plus 5 ? 2x – 5, with the restriction x ≠ –five over 2 2x + 5, with the restriction x ≠ five over 2 2x – 5, with the restriction x ≠ five over 2 2x + 5, with the restriction x ≠ –five over 2

OpenStudy (mathstudent55):

This? \( \dfrac{4x^2 - 25}{2x + 5} \)

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

Do you know how to factor a difference of two squares?

OpenStudy (mathstudent55):

This is how you factor the difference of two squares: \( a^2 - b^2 = (a + b)(a - b) \)

OpenStudy (anonymous):

...

OpenStudy (mathstudent55):

In the numerator, you have \(4x^2 - 25 \) \(4x^2\) is the square of \(2x\) \(25\) is the square of \(5\) Your numerator is the difference of two squares.

OpenStudy (mathstudent55):

That means you can factor the nmumerator as: \( \dfrac{4x^2 - 25}{2x + 5} = \dfrac{(2x + 5)(2x - 5)}{2x + 5} \)

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

Now you can divide the numerator and denominator by the common factor of 2x + 5: \( \ = \dfrac{(2x + 5)(2x - 5)}{2x + 5} \) \( \ = 2x - 5 \)

OpenStudy (mathstudent55):

The fraction simplifies to only 2x - 5, but since the expression started with a denominator, you need to include the restrictioon that the denominator contains. The denominator can't equal zero. To find what value of x makes the denominator zero, set the original denominator equal to zero and solve for x: \( 2x + 5 = 0\) \( 2x = -5\) \(x = -\dfrac{5}{2} \) Finally you have the complete answer: \(2x - 5\), \(x \ne -\dfrac{5}{2}\)